OpenStudy (anonymous):

Still not getting it: x-4y=9 2x-3y=11

OpenStudy (anonymous):

thats too bad

OpenStudy (anonymous):

Why do you have to go and be a jerk? that uncalled for

OpenStudy (anonymous):

x-4y=9 2x-3y=11 2x-8y=18 2x-3y=11 -5y=7 y=-5/7 x=9+4(-5/7)

jimthompson5910 (jim_thompson5910):

Solve x-4y=9 for x x-4y=9 x=9+4y --------------------- Now move onto the second equation 2x-3y=11 and plug in x = 9+4y (what we just solved for above) 2(9+4y)-3y=11 Now solve for y 2(9+4y)-3y=11 18+8y-3y=11 5y+18=11 5y=11-18 5y=-7 y=-7/5 Now go back to x=9+4y and use y=-7/5 to find x x=9+4y x=9+4(-7/5) x=9-28/5 x=45/5-28/5 x=17/5 So the solutions are x=17/5 and y=-7/5

OpenStudy (anonymous):

multiply 1st equation by 2 2(x-4y=9)------>2x-8y=18 subtract 2nd equation from 1st 2x-8y=18 -2x+3y=-11 ------------ -5y=7 y=-7/5 now plug y into one of the equation x-4(-7/5)=9 x+(28/5)=9 x=9-(28/5) x=(45-28)/5 x=17/5 therefore,x=17/5 and y=-7/5